A car with a mass of 1000 kg is moving directly East at a rate of 10 m/s collides with a 1500 kg car moving directly North at 18 m/s. If this is a perfectly inelastic collision, what is the resulting momentum and direction of the final two car mass? What is the masses velocity?


1
Expert's answer
2021-06-07T15:01:51-0400

Given:

"v_1=10\:\rm m/s"

"v_2=18\:\rm m/s"

"m_1=1000\:\rm kg"

"m_2=1500\:\rm kg"

The law of conservation of momentum says


"m_1v_1=p_x\\
m_2v_2=p_y"

The magnitude of final momentum


"p=\sqrt{p_x^2+p_y^2}=\sqrt{(m_1v_1)^2+(m_2v_2)^2}"

"=\sqrt{(1000\times 10)^2\!+\!(1500\times 18)^2}=28792\: \rm kg\cdot m/s"

Direction


"\theta=\tan^{-1}\frac{p_y}{p_x}=\tan^{-1}\frac{1500\times 18}{1000\times 10}=70^{\circ} \rm N\: of\: E"

The velosity


"v=\frac{p}{m_1+m_2}=\frac{28792}{1000+1500}=11.5\:\rm m/s"